Compiler-directed Type Reconstruction for Polymorphic

Languages

Computation Structures Group Memo 348 March 29, 1993 Shail Aditya Alejandro Caro

In Proceedings of the Conference on Functional Programming Languages and Computer Architecture, Copenhagen, Denmark, June 9-11, 1993

The research described in this paper was funded in part by the Advanced Research Projects Agency of the Department of Defense under Office of Naval Research contract N00014-89-J-1988.

Compiler-directed Type Reconstruction for Polymorphic Languages

Shail Aditya Alejandro Caro

MIT Laboratory for Computer Science 545 Technology Square, Cambridge, MA 02139

  {shail,acaro}@abp.lcs.mit.edu

March 29, 1993